A cube is 8inches on each side. What is the height of a cylinder having the same volume, if its radius is 4 inches? Round to nearest tenth.
Answer by mananth(16949)

the side of the cube =8 inches
volume of cube = 8^3=512
volume of cylinder = 3.14 * 4^2*h
512= 3.14*16*h
512/ 3.14*16=h
h=10.19